# Semantic Component Analysis: Discovering Patterns in Short Texts Beyond Topics

### v2

@@ Abstract
Topic modeling is a key method in text analysis, but existing approaches are limited by assuming one topic per document or fail to scale efficiently for large, noisy datasets of short texts. We introduce Semantic Component Analysis (SCA), a novel topic modeling technique that overcomes these limitations by discovering multiple, nuanced semantic components beyond a single topic in short texts which we accomplish by introducing a decomposition step to the clustering-based topic modeling framework. We evaluate SCA on Twitter datasets in English, Hausa and Chinese. It achieves competetive coherence and diversity compared to BERTopic, while uncovering at least double the semantic components and maintaining a noise rate close to zero. Furthermore, SCA is scalable and effective across languages, including an underrepresented one.
Semantic Component Analysis:
Discovering Patterns in Short Texts Beyond Topics
@@ 1 Introduction
Topic modeling is an essential tool for automated text analysis in digital humanities, computational social science, and related disciplines Grimmer et al. (2022); Rauchfleisch et al. (2023); Blei (2012); Maurer and Nuernbergk (2024). It aims to find high-level semantic patterns, the topics, in a corpus and then assign them to documents. The state-of-the-art for large-scale datasets, BERTopic Grootendorst (2022), assumes a single topic per document, which often does not hold even for short texts Paul and Girju (2010), as can be seen in Figure 1. BERTopic and related approaches thus do not model a large portion of the information in the documents. Additionally, while they achieve high topic coherence and diversity (defined in Appendix B.2), a high rate of samples is usually assigned to no topic, i.e. they have a high noise rate Egger and Yu (2022).
@@
Figure 1: A Tweet from the Trump dataset (Brown, 2021) decomposed into its semantic components as defined by the Top-3 component activation scores of iteration 1 (components 6, 24, and 0) and subsequent iterations (components 56, 118, 70). Component 0 refers to the most dominant topic in the Trump dataset, of which a majority of Tweets is about Donald Trump himself.
@@
To overcome these limitations, we propose Semantic Component Analysis (SCA) as a novel method to discover multiple patterns per document, which we coin semantic components. SCA is a topic modeling technique, that — in addition to topics — formalizes the discovery of nuanced semantic components. It scales to large datasets while maintaining a noise rate close to zero and matches BERTopic in topic coherence and diversity. Based on clustering sequence embeddings as proposed by Angelov (2020), for each cluster, we introduce a decomposition step and iterate over the residual embeddings (see Figure 2). This decomposition enables us to represent a sample’s content independently of the semantic components found in previous iterations and to find additional components.
@@
Figure 2: Illustration of our proposed method, that clusters residual embeddings obtained from linear decompositions of the embedding space. Top-Right: Top-4 components (by cluster size) of the first iteration equivalent to a BERTopic run. Bottom-right: components uncovered in subsequent iterations (see Appendix C.2 for more).
@@
We evaluate our method on three Twitter datasets, one of Tweets by Donald Trump (Brown, 2021), another one of Tweets in Hausa (Muhammad et al., 2022), and finally one compiled from Chinese-language news outlet accounts that we scraped ourselves and publish alongside this paper. We show that our method is able to match the performance of BERTopic in terms of topic coherence and diversity, while additionally providing a large number of semantic components missed by the state-of-the-art approach. We also show that our method can provide an interpretable transformation that matches the performance of the base embedding in a variety of downstream tasks.
Our main contributions are as follows:
  • We propose a novel topic modeling method for short texts, Semantic Component Analysis (SCA), that can discover multiple patterns per sample, which we coin "semantic components".
  • We show that our method is able to match state-of-the-art topic modeling, while providing additional, more nuanced insights into the data and achieving a noise rate close to zero.
  • We publish a dataset of Tweets from Chinese news outlets and use a dataset of Tweets in Hausa, a low-resource language, to show the generalizability of our method beyond English.
@@ 2 Related Work
Topic modeling’s goal is to find semantic patterns in a corpus and to assign them to the documents. Latent Dirichlet Allocation (LDA, Blei et al. 2003; Pritchard et al. 2000) works well on small datasets of long texts and has been generalized to larger datasets Hoffman et al. (2012). Modeling a distribution over topics for each document, it can account for multi-topic documents. However, it performs poorly on short texts Ushio et al. (2021); de Groot et al. (2022), is generally vulnerable to noisy data, and produces topics that lack domain adaption Egger and Yu (2022). Recent advancements have integrated neural networks (Srivastava and Sutton, 2017; Liu et al., 2015; Dieng et al., 2019; Qiang et al., 2016; Zhao et al., 2021). Topic modeling based on LLMs Pham et al. (2024); Doi et al. (2024) has also gained attention, though their cost is a barrier for most large-scale applications. Methods based on clustering embeddings stand out because of their efficiency, domain-adaption, and topic coherence on large corpora of short texts. Top2Vec (Angelov, 2020) clusters document embeddings generated by Doc2Vec (Le and Mikolov, 2014). There, UMAP (McInnes et al., 2020) reduces embedding dimensionality mitigating the "curse of dimensionality" (Reimers and Gurevych, 2021), and clustering is performed using HDBSCAN (Campello et al., 2015). BERTopic (Grootendorst, 2022) improves the Top2Vec framework by using SBERT embeddings (Reimers and Gurevych, 2019) and c-TF-IDF weighting for representing the topics. It is considered state-of-the-art for large datasets of short texts Egger and Yu (2022). However, BERTopic does not provide a way to generalize to multi-topic documents. The cluster-hierarchy of HDBSCAN could be used to assign multiple topics per sample, but this approach is insufficient for finding patterns that are not subtopics of each other. Additionally, BERTopic’s results usually suffer from a high noise rate Egger and Yu (2022), which can range at up to 74% de Groot et al. (2022). In summary, the topic modeling field does not provide a solution for scenarios with large, noisy datasets of short texts containing multiple patterns of interest per sample, that matches the state-of-the-art topic quality.
@@ 3 Methodology
SCA is a clustering pipeline that introduces a decomposition step to find semantic components in the embedding space, which are represented by unit vectors. For clustering, we employ the cosine distance in UMAP to map the embeddings to a 5-dimensional, Euclidean space on which HDBSCAN operates, mirroring the setup in BERTopic. Visualized in Figure 2, it works as follows:
@@
    $\mathbf{v}_{i}=\frac{\mathbf{v}^{\prime}_{i}}{||\mathbf{v}^{\prime}_{i}||},\ \ \mathbf{v}^{\prime}_{i}=\frac{1}{|C_{i}|}\sum_{\mathbf{x}_{j}\in C_{i}}\mathbf{x}_{j}$ (1)
@@
    Decompose the $\mathbf{x}_{j}$ into their linear components along each $\mathbf{v}_{i}$ and the residual $\mathbf{x}_{i}^{\prime}$:
@@
    where $\alpha_{i,j}=\frac{\langle\mathbf{x}_{j},\mathbf{v}_{i}\rangle}{||\mathbf{x}_{j}||}$
@@
    Represent and Merge components by c-TF-IDF weighting and token overlap
After each decomposition step in 3, we set $\mathbf{x}_{j}=\mathbf{x}_{j}^{\prime}$ for all $j$ and repeat it for the next cluster. The algorithm converges when the residuals are zero. We introduce other practical stopping criteria listed in Appendix A.3. Note, that SCA has two hyperparameters: $\mu\in[0,1]$, controlling the amount of the decomposition, and a threshold $\alpha$, which is compared against $\alpha_{i,j}$, the cosine similarity between $\mathbf{x}_{j}$ and $\mathbf{v}_{i}$. See Appendix A.1 for a detailed discussion of $\alpha$ and $\mu$. We assume that each cluster found represents a semantic component. To obtain an interpretable representation, we follow BERTopic and use c-TF-IDF to rank the tokens according to their importance for a cluster storing the Top-10 set of tokens for each cluster.
We introduce a merging step that computes the token overlap between two components as the fraction of tokens their representations share. Here, we introduce a threshold $\theta\in[0,1]$ and merge two components into a multi-dimensional component if the token overlap is larger than $\theta$. We keep the representation of the component that appeared earlier in the procedure. That way, we avoid duplicates arising in larger clusters spread across multiple dimensions of the embedding space.
Finally, we include the medoid of each cluster, which is the sample closest to the centroid in terms of cosine similarity, as a representation.
We note, that SCA employs the same modular approach as BERTopic, which allows the embedding, dimensionality reduction, clustering, and representation method to be replaced by similar methods.111see https://maartengr.github.io/BERTopic/ Furthermore, our specific choices for these parts of the model, namely UMAP and HDBSCAN, make the first iteration of SCA equivalent to a BERTopic run, which allows for a direct comparison of the methods.
@@ 4 Data & Evaluation
We evaluate our method on three datasets of different sizes and languages.222code and data provided in the supplementary materials For an English-language use-case, we include all Tweets by Donald Trump until December 2024 (60K samples, Brown 2021). To cover a low-resource language, we include a set of 512K Tweets in Hausa collected by Muhammad et al. (2022). To evaluate on a high-resource non-English language, we collect and publish 1.5M Chinese-language Tweets from a list of news outlets (details in Appendix B.1). We evaluate our method in terms of number of components, noise rate, topic coherence, and diversity (see Appendix B.2 for definitions) and compare it to the highly popular BERTopic. Furthermore, we perform a grid run over the hyperparameters to assess their influence. To check if the components found are indeed beyond what is found by BERTopic, we compute the token overlap between the components found by SCA and the topics found by BERTopic, as well as the sample overlap with the cluster hierarchy in HDBSCAN of BERTopic and the components found by SCA (see Appendix B.2 for details). Finally, we evaluate the SCA’s transformation into the component space (see Appendix A.2 for details) on the MTEB benchmark (Muennighoff et al., 2023), which consists of various downstream tasks, including sequence classification, summarization and retrieval. We do this to quantify the faithfulness of the representation provided by SCA-components and compare against the base embedding model and a PCA.
@@ 5 Results
Statistic Trump Hausa Ch. News
BERTopic SCA BERTopic SCA BERTopic SCA
Number of Components $\uparrow$ 55 182 102 331 212 594
Noise Rate $\downarrow$ 0.275 0.000 0.574 0.001 0.429 0.000
NPMI Coherence $\uparrow$ -0.135 -0.168 -0.028 -0.061 0.186 0.120
CV Coherence $\uparrow$ 0.370 0.360 0.431 0.400 0.683 0.595
Topic Div. $\uparrow$ 0.947 0.703 0.909 0.804 0.929 0.803
Avg. Max. Sample Overlap $\in[0,1]$downarrow 0.079 0.093 0.077
Avg. Max. Token Overlap $\in[0,10]$downarrow 1.387 1.852 1.318
[…]
@@
Besides the qualitative results in Figure 2, we provide statistical results in Table 1. For all datasets, the method finds a large number of components, with the Trump dataset yielding 182 components in total after merging, the Hausa dataset 331 and the Chinese News dataset 594, surpassing the 55, 102, and 212 respective topics found by BERTopic. These components relate to patterns distinct from the topics found by BERTopic, which is evident from the low average maximum token overlap scores of 1.85, 1.32 and 1.39 on the three datasets respectively. On the Trump dataset, only 8 components with a token overlap of higher than 0.5 are merged, on the Hausa dataset it is 9, and for the Chinese News dataset 19. The average maximum sample overlaps with any of BERTopic’s hierarchical subtopics are small and between 0.077 and 0.093 on all datasets. This indicates that the additional components are not sub- or supertopics of the topics found by BERTopic. The noise rate is close to zero for all datasets, i.e. almost all samples are assigned at least one component. The topic coherence scores are high for all datasets, with the Chinese dataset showing the highest NPMI coherence of 0.120 and the Trump dataset showing the lowest of -0.168, all matching the baseline. The topic diversity is high for all datasets, between 0.7 and 0.8, but about 10-15% lower than in the baseline.
@@
In sum, this shows that the quality of the components found matches that of BERTopic while providing a more nuanced view of the data through the large number of extra components found and the low noise rate. Comparing the results over the hyperparameters on the Trump dataset, we find that for $\alpha>0.7$ and $\mu<0.4$, the method reliably finds a large number of semantic components at least double the number of topics found by BERTopic, usually ranging from three to four times.
@@
The additional components found relate to patterns that intuitively match the expected distribution of the datasets (full Top-10 component tables in Appendix C). For example, in the Trump dataset, beyond the topics about Trump himself, entrepreneurship, and election campaigning found by BERTopic, we uncover components about "fake news", inspiration and role models, paying condolences, and the "witch hunt" narrative among others (see Appendix C.2) underlining SCA’s ability to find important patterns not captured by the baseline. This view is reinforced by the component distributions of single samples, as shown in Figure 1. Similarly, in the Chinese dataset, the additional components relate to important motives in Chinese news reporting, such as protests, family, the trade war with the US, and currency (see Appendix C.3). In the Hausa dataset, the additional insights include components about poverty, courts and law, and work (see Appendix C.4). However, we note, that in this low-resource scenario, some components relate to semantically more narrow concepts, which is possibly due to a lack of expressive power in the base embeddings.
@@
Finally, the results of the topic distribution provided by SCA can match the performance of PCA and the base embedding model in the MTEB benchmark (see Appendix C.5), indicating the method’s faithfulness in representing the information of the samples in the component space.

### v3

# Semantic Component Analysis: Introducing Multi-Topic Distributions to Clustering-Based Topic Modeling

@@ Abstract
Topic modeling is a key method in text analysis, but existing approaches fail to efficiently scale to large datasets or are limited by assuming one topic per document. Overcoming these limitations, we introduce Semantic Component Analysis (SCA), a topic modeling technique that discovers multiple topics per sample by introducing a decomposition step to the clustering-based topic modeling framework. We evaluate SCA on Twitter datasets in English, Hausa and Chinese. There, it achieves competetive coherence and diversity compared to BERTopic, while uncovering at least double the topics and maintaining a noise rate close to zero. We also find that SCA outperforms the LLM-based TopicGPT in scenarios with similar compute budgets. SCA thus provides an effective and efficient approach for topic modeling of large datasets.
Semantic Component Analysis: Introducing Multi-Topic Distributions to Clustering-Based Topic Modeling
@@ 1 Introduction
Topic modeling is an essential tool for automated text analysis in digital humanities, computational social science, and related disciplines Grimmer et al. (2022); Rauchfleisch et al. (2023); Blei (2012); Maurer and Nuernbergk (2024). It aims to find high-level semantic patterns, the topics, in a corpus and then assign them to documents. The state-of-the-art (SOTA) for large-scale datasets, BERTopic Grootendorst (2022), assumes a single topic per document, which often does not hold even for short texts Paul and Girju (2010), as can be seen in Figure 1. Additionally, a high rate of samples is usually assigned to no topic, i.e. they have a high noise rate Egger and Yu (2022). LLM-based approaches like TopicGPT Pham et al. (2024) perform very well, but due to their high computational cost, it is questionable how well they scale to large datasets under resource constraints.
@@
Figure 1: A Tweet from the Trump dataset (Brown, 2021) decomposed into its semantic components as defined by the TOP3 component activation scores of iteration 1 (components 6, 24, and 0) and subsequent iterations (components 56, 118, 70).
@@
To overcome these limitations, we propose Semantic Component Analysis (SCA) as a novel method to discover multiple topics per document. SCA is a topic modeling technique, that formalizes the discovery of semantic components, which intuitively are mathematical directions in the embedding space, each corresponding to a topic. It scales to large datasets while maintaining a noise rate close to zero and matches BERTopic in topic coherence and diversity. Based on clustering sequence embeddings as proposed by Angelov (2020), for each cluster, we introduce a decomposition step and iterate over the residual embeddings (see Figure 2). This decomposition enables us to represent a sample’s content independently of the topics found in previous iterations and to find additional topics in the residuals.
@@
Figure 2: Illustration of our proposed method, that clusters residual embeddings obtained from linear decompositions of the embedding space. Top-Right: TOP4 components (by cluster size) of the first iteration equivalent to a BERTopic run. Bottom-right: components uncovered in subsequent iterations (see Appendix C.2 for more).
@@
We evaluate our method on four datasets, one of Tweets by Donald Trump (Brown, 2021), another one of Tweets in Hausa (Muhammad et al., 2022), one of Tweets by Chinese-language news outlet accounts that we publish alongside this paper, and a collection of bills by the US congress Adler and Wilkerson (2012). We show that our method is able to match BERTopic in terms of topic coherence and diversity, while additionally providing a large number of topics missed by the SOTA approach. Furthermore, we show that given a comparable compute budget, SCA outperforms the LLM-based TopicGPT baseline on the Bills dataset, a benchmark proposed by Pham et al. (2024). Lastly, we also show that our method provides an interpretable transformation matching the base embedding in a variety of downstream tasks.
@@ 2 Related Work
Topic modeling’s goal is to find semantic patterns in a corpus and to assign them to the documents. Latent Dirichlet Allocation (LDA, Blei et al. 2003; Pritchard et al. 2000) models multi-topic documents, works well on small datasets of long texts, and has been generalized to larger datasets Hoffman et al. (2012). However, it performs poorly on short texts Ushio et al. (2021); de Groot et al. (2022), is generally vulnerable to noisy data, and lacks domain adaption Egger and Yu (2022). Recent advancements have integrated neural networks (Srivastava and Sutton, 2017; Liu et al., 2015; Dieng et al., 2019; Qiang et al., 2016; Zhao et al., 2021), but they tend to struggle with stability and alignment with human-determined categories Hoyle et al. (2022). Topic modeling based on LLMs Pham et al. (2024); Doi et al. (2024) is promising, but their cost is a barrier for large-scale applications (see discussion in Appendix B.2) and their generated topics tend to be too generic or general Li et al. (2025). Methods based on clustering embeddings stand out because of their efficiency, domain-adaption, and topic coherence on large corpora of short texts. Top2Vec (Angelov, 2020) clusters document embeddings generated by Doc2Vec (Le and Mikolov, 2014). There, UMAP (McInnes et al., 2020) reduces embedding dimensionality mitigating the "curse of dimensionality" (Reimers and Gurevych, 2021), and clustering is performed using HDBSCAN (Campello et al., 2015). BERTopic (Grootendorst, 2022) improves the Top2Vec framework by using SBERT embeddings (Reimers and Gurevych, 2019) and c-TF-IDF weighting for representing the topics. It is considered SOTA for large datasets of short texts Egger and Yu (2022). However, BERTopic does not provide a way to generalize to multi-topic documents. The cluster-hierarchy of HDBSCAN could be used to assign multiple topics per sample, but this approach is insufficient for finding patterns that are not subtopics of each other. Additionally, BERTopic’s results usually suffer from a high noise rate Egger and Yu (2022), which can reach up to 74% de Groot et al. (2022). In summary, the topic modeling field does not provide a solution for scenarios with large, noisy datasets of short texts containing multiple topics per sample, that matches the SOTA topic quality.
@@ 3 Methodology
SCA is a clustering pipeline that introduces a decomposition step to find unit vectors in the embedding space—which we call semantic components—that each correspond to a topic. For clustering, we employ the cosine distance in UMAP to map the embeddings to a 5-dimensional, Euclidean space on which HDBSCAN operates, mirroring the setup in BERTopic. Visualized in Figure 2, it works as follows:
@@
[…]
@@
[…]
@@
    Represent components by c-TF-IDF
  • 6.
    Merge topics whose token representations overlap by more than a threshold $\theta$.
After each decomposition step in 3, we set $\mathbf{x}_{j}=\mathbf{x}_{j}^{\prime}$ for all $j$ and repeat it for the next cluster. The algorithm converges when the residuals are zero. We introduce other practical stopping criteria listed in Appendix A.3. Note that SCA has two hyperparameters: $\mu\in[0,1]$, controlling the amount of the decomposition, and a threshold $\alpha$, which is compared against $\alpha_{i,j}$, the cosine similarity between $\mathbf{x}_{j}$ and $\mathbf{v}_{i}$. See Appendix A.1 for a detailed discussion of $\alpha$ and $\mu$. We assume that each cluster found represents a topic. To obtain an interpretable representation, we follow BERTopic and use c-TF-IDF to rank the tokens according to their importance for a cluster storing the TOP10 set of tokens for each cluster.
Merging step.
To avoid duplicates arising in larger clusters spread across multiple dimensions, we introduce a merging step. First, for each pair of topics, we compute the token overlap $O(R_{1},R_{2}):=\frac{1}{10}|R_{1}\cap R_{2}|$ between their token representations $R_{1},R_{2}$, where the division by $10$ serves as a normalization factor for the number of tokens per representation. If $O(R_{1},R_{2})>\theta$ for some threshold $\theta\in[0,1]$, we merge the two corresponding topics by assigning all the same topic ID and keeping the token and component representations of the topic that appeared earlier.
Topic assignment.
SCA can assign topics in two ways: (i) Assign topics based on the clusterings in each iteration or (ii) obtain a full topic distribution by mapping the residual embeddings $\mathbf{x}^{{}^{\prime}}_{j}$ to the component space, where each dimension $\langle\mathbf{x}^{{}^{\prime}}_{j},\mathbf{v}_{i}\rangle$ is interpreted as relevance of the topic represented by component $\mathbf{v}_{i}$ (see Appendix A.2).
SCA has the same modular approach as BERTopic, in which the embedding, dimensionality reduction, clustering, and representation method can be replaced.111see https://maartengr.github.io/BERTopic/ Our specific choices for these modules, namely SBERT, UMAP, HDBSCAN, and c-Tf-IDF, make the first iteration of SCA equivalent to a BERTopic run, which allows for a direct comparison of the methods.
@@ 4 Data & Evaluation
Datasets.
We evaluate our method on four datasets of different sizes and languages. [^2] https://github.com/mainlp/semantic_components For an English-language use-case, we include all Tweets by Donald Trump until December 2024 (60K samples, Brown 2021). To cover a low-resource language, we include a set of 512K Tweets in Hausa collected by Muhammad et al. (2022). To evaluate on a high-resource non-English language, we collect and publish 1.5M Chinese-language Tweets from a list of news outlets (details in Appendix B.1). To evaluate in another domain against a ground truth, we use the Bills dataset (32K samples, Adler and Wilkerson 2012) used by Pham et al. (2024).
Evaluation metrics.
We evaluate our method in terms of number of topics, noise rate, topic coherence, and diversity (see Appendix B.3 for definitions) and compare it to LDA and BERTopic on the datasets without ground truth. Furthermore, we perform a grid run over the hyperparameters to assess their influence. To check if the topics found are indeed beyond what is found by BERTopic, we compute the token overlap between the topics found by SCA and the topics found by BERTopic, as well as the sample overlap with the cluster hierarchy in HDBSCAN of BERTopic and the topics found by SCA (see Appendix B.3 for details). Evaluating on the Bills dataset, we compare SCA to TopicGPT Pham et al. (2024) in terms of Adjusted Rand Index (ARI), Normalized Mutual Information (NMI), and Purity ($P_{1}$) as proposed by them. Note that running TopicGPT requires a high budget of several thousand dollars for large datasets (see Discussion in Appendix B.2). We thus also compare to a version of TopicGPT with similar compute budget as SCA.
MTEB evaluation.
Finally, we evaluate the SCA’s transformation into the component space (see Appendix A.2 for details) on the MTEB benchmark (Muennighoff et al., 2023), which consists of various downstream tasks, including sequence classification, summarization and retrieval. We do this to quantify the faithfulness of the representation provided by SCA-components and compare against the base embedding model and a PCA.
Trump Hausa Chinese News
BT SCA LDA BT SCA LDA BT SCA LDA
#Topics $\uparrow$ 55 182 182 102 331 331 212 594 594
Noise R. $\downarrow$ 0.275 0.000 0.003 0.574 0.001 0.05 0.429 0.000 0.802
NPMI $\uparrow$ -0.135 -0.168 -0.139 -0.028 -0.061 -0.156 0.186 0.120 -0.149
CV $\uparrow$ 0.370 0.360 0.334 0.431 0.400 0.393 0.683 0.595 0.198
Topic D. $\uparrow$ 0.947 0.703 0.698 0.909 0.804 0.444 0.929 0.803 0.003
Bills
BT SCA LDA TGPT TGPT TGPT
#Topics 79 79 79 23 15 79
Param. 0.3B 0.3B n/a 1B 8B unk.
$P_{1}$ $\uparrow$ 0.42 0.42 0.39 0.13 0.16 0.57
ARI $\uparrow$ 0.16 0.18 0.21 0.00 0.00 0.42
NMI $\uparrow$ 0.38 0.38 0.47 0.01 0.09 0.52
Table 1: Topic modeling statistics. BERTopic and SCA use the same hyperparameters. Arrows ($\uparrow\downarrow$) indicate optimal metric directions (see Appendix B.3). Results marked with are copied from Pham et al. (2024), who report using GPT-3.5 through the OpenAI API. Additional runs of TopicGPT use Llama-3.2-1B-Instruct and Llama-3.2-8B-Instruct Grattafiori (2024). We follow their evaluation to compare SCA to TopicGPT, which is restricted to the TOP79 topics to ensure comparability with the LLM based approach. Since the two TopicGPT baselines do not find this many topics, we report the results for the respective smaller numbers of topics.
@@ 5 Results
Statistical results.
@@
We provide statistical results in Table 1. For all datasets, SCA finds a large number of topics, with the Trump dataset yielding 182 topics in total after merging, the Hausa dataset 331 and the Chinese News dataset 594, surpassing the 55, 102, and 212 respective topics found by BERTopic. These topics relate to patterns distinct from the topics found by BERTopic, which is evident from the low average maximum token overlap scores of 1.85, 1.32 and 1.39 on the three datasets respectively. On the Trump dataset, only 8 topics with a token overlap of higher than 0.5 are merged, on the Hausa dataset it is 9, and for the Chinese News dataset 19. The average maximum sample overlaps with any of BERTopic’s hierarchical subtopics are small and between 0.077 and 0.093 on all datasets. This indicates that the additional topics are not sub- or supertopics of the topics found by BERTopic.
The noise rate is close to zero for all datasets, i.e. almost all samples are assigned at least one topic. The topic coherence scores are high for all datasets, with the Chinese dataset showing the highest NPMI coherence of 0.120 and the Trump dataset showing the lowest of -0.168, all matching the baseline. The topic diversity is high for all datasets, between 0.7 and 0.8, but about 10-15% lower than in the baseline.
@@
In sum, this shows that the quality of the topics found matches that of BERTopic while providing a more nuanced view of the data through the large number of extra topics found and the low noise rate. Ablating the hyperparameters on the Trump dataset, we find that for $\alpha>0.7$ and $\mu<0.4$, SCA reliably finds at least double the number of topics found by BERTopic, usually ranging up to four times.
Tweet 1: ALWAYS BORROW MONEY FROM A PESSIMIST BECAUSE HE WILL NEVER EXPECT IT TO BE PAID BACK!
TOP1: entrepreneurs, yourself, success, think, focus, champion, learn, art, passion, touch
TOP2: (2+ It.) charity, money, ads, million, bankrupt, bankruptcy, website, ad, raised, donate
TOP3: impeachment, collusion, democrats, investigation, hoax, whistleblower, evidence, hunt, dems, witch
Tweet 2: "Leverage: don’t make deals without it.” – The Art of the Deal
TOP1: (1st It.) entrepreneurs, yourself, success, think, focus, champion, learn, art, passion, touch
TOP2: (2+ It.) art, deal, negotiation, desperate, seem, possibly, deals, tip, worst, theartofthedeal
TOP3: (1st It.) impeachment, collusion, democrats, investigation, hoax, whistleblower, evidence, hunt, dems, witch
Tweet 3: 95% Approval Rating in the Republican Party. Thank you!
TOP1: (2+ It.) approval, rating, overall, party, republican, 52, 51, 53, 93, 50
TOP2: (1st It.) impeachment, collusion, democrats, investigation, hoax, whistleblower, evidence, hunt, dems, witch
TOP3: (1st It.) poll, leads, surges, lead, bush, surging, postdebate, tops, 28, favorability
Tweet 4: Our govt is so pathetic that some of the billions being wasted in Afghanistan are ending up with terrorists
TOP1: (1st It.) impeachment, collusion, democrats, investigation, hoax, whistleblower, evidence, hunt, dems, witch
TOP2: (1st It.) isis, oil, fighters, caliphate, attack, chemical, rebels, terrorist, troops, terrorists
TOP3: (2+ It.) leadership, leader, worst, incompetent, need, politicians, needs, stamina, leaders, lead
Table 2: Samples from the Trump dataset for qualitative error analysis. Note that the topics are taken from a different run than the one described in the Appendix, but with the same hyperparameters. Top two samples were selected at random from the entrepreneurship topic and the bottom two were selected from the noise cluster of iteration 1, i.e. would have been classified as noise by BERTopic. We also highlight the largest topic colored in red. Topics shown are the TOP3 topics assigned by the scores of the transformation into the component space and annotated on whether they were discovered in the first iteration or subsequent iterations.
Exploration of additional topics.
@@
The additional topics found intuitively match the expected distribution of the datasets (full TOP10 topic tables in Appendix C). For example, in the Trump dataset, beyond the topics found by BERTopic (about Trump himself, entrepreneurship, election campaigning), we uncover topics about "fake news", inspiration and role models, paying condolences, and the "witch hunt" narrative among others (see Appendix C.2) underlining SCA’s ability to find important topics not captured by the baseline.
This view is reinforced by the topic distributions of single samples, as shown in Figure 1. Similarly, in the Chinese dataset, the additional topics relate to important motives in Chinese news reporting, such as protests, family, the trade war with the US, and currency (see Appendix C.4). In the Hausa dataset, the additional insights include topics about poverty, courts and law, and work (see Appendix C.5). In this low-resource scenario, some topics relate to semantically more narrow concepts, which might be caused by lack of expressive power in the embeddings.
Error analysis.
While the discovered topics are generally high quality, we observe several recurring error patterns. In all datasets, SCA produces a power-law distribution of topic sizes, with a few large and many smaller topics. The largest topics tend to be generic and over-assigned, as seen in Table 2. This is likely due to HDBSCAN, which tends to connect clusters in dense areas of the embedding space. However, in later iterations SCA refines the broad entrepreneurship topic into more specific topics such as money and negotiation assigned to Tweets 1 and 2 respectively. We also sample two Tweets from the iteration-1 noise cluster—i.e., Tweets BERTopic failed to assign—and find that SCA correctly identifies topics like approval ratings, polling, terrorism, and leadership criticism. We also note that the largest topic, which seems to encompass patterns such as impeachment and the Democratic party, is wrongly predicted across all samples. Lastly, while most TOP10 token lists support interpretation, some topics, especially the generic ones, are hard to interpret.
Comparison with TopicGPT.
Comparing against TopicGPT, we find that both SCA and BERTopic achieve similar performance in terms of purity, NMI and ARI outperforming the TopicGPT baselines based on the Llama models (i.e. with similar compute budgets). The good results reported by Pham et al. (2024) are based on GPT-3.5, suggesting that compute-intensive and expensive models are necessary for TopicGPT to perform well; which might not be available in application settings like social science.
MTEB evaluation of transformed embeddings.
@@
Finally, the results of the topic distribution provided by SCA can match the performance of PCA and the base embedding model in the MTEB benchmark (see Appendix C.6), indicating the method’s faithfulness in representing the information of the samples in the component space. This suggests that the transformed embeddings provided by SCA can be seen as an interpretable alternative to the base embedding model at no cost in performance.